# Feshbach-Villars Formalism for a Spin-1/2 Particle in Curved Spacetime

### v1

@@ III Feshbash-Villars of a particle of spin-1/2 in curved spacetime: The formalism
III.2.1 Special Case: Uniform Fields
  • Uniform Magnetic Field $(\vec{E}=0)$
    $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=-e\vec{\Sigma}\cdot\vec{B}.$ (45)
  • Uniform Electric Field $(\vec{B}=0)$
    $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=e\gamma^{0}\vec{\gamma}\cdot\vec{E}.$ (46)
IV Conclusion
The Feshbach-Villars formalism for a spin-1/2 particle in curved spacetime provides a robust method for exploring relativistic quantum mechanics in the presence of gravitational and electromagnetic fields. By deriving the Hamiltonian form of the Dirac equation and applying the FV transformation, we have developed a two-component formulation that clearly separates positive and negative energy states. This formalism presents several advantages:
  • It allows a more intuitive interpretation of the wavefunction components with respect to particle and antiparticle contributions.
  • The resulting Hamiltonian, expressed in terms of Pauli matrices, facilitates the analysis of spin-dependent effects in curved spacetime.
  • The approach naturally incorporates both gravitational and electromagnetic interactions, making it suitable for exploring various physical scenarios.
Our analysis in both (1+2) and (1+3) dimensions showcases the versatility of the formalism and highlights the specific forms of spin-field interactions across different spacetime configurations. The explicit treatment of the term involving the electromagnetic field tensor and spin matrices provides a clear understanding of how spin couples to external fields in curved spacetime. This work lays the groundwork for future studies on relativistic quantum effects in strong gravitational fields, the behavior of spin in curved spacetime, and potential applications in areas such as early universe cosmology and the physics of compact objects. Further research could investigate the implications of this formalism for quantum field theory in curved spacetime and explore possible experimental tests of the predicted effects.
?refname?
  • 1 (2025) H. Feshbach and F. M. H. Villars, Rev. Modern Phys. 30, 24 (1958).
  • 2 (2025) B. A. Robson and D. S. Staudte, J. Phys. A : Math. Gen 29, 157 (1996).
  • 3 (2025) D. S. Staudte, J. Phys. A 29, 169 (1996).
  • 4 (2025) M. Merad, L. Chetouani, and A. Bounames, Phys. Lett. A 267, 225 (2000).
  • 5 (2025) A. Bounames and L. Chetouani, Phys. Lett. A 279, 139 (2001).
  • 6 (2025) S. Haouat and L. Chetouani, Eur. Phys. J. C 41, 297 (2005).
  • 7 (2025) B Guettou1 and L Chetouani, Phys. Scr. 74, 12–18 (2006).
  • 8 (2025) N. Brown, Z. Papp, and R. M. Woodhouse, Few-Body Systems 57, 103 (2015).
  • 9 (2025) B. Motamedi, T. Shannon, and Z. Papp, Few-Body Systems 60, 65 (2019).
  • 10 (2025) D. Wingard, A. Garcia Vallejo and Z. Papp, Few-Body Syst 65, 30 (2024).
  • 11 (2025) A. J. Silenko, Phys. Rev. A 77, 012116 (2008).
  • 12 (2025) A. J. Silenko, Phys. Rev. D 88, 045004 (2013).
  • 13 (2025) A. Bouzenada, A. Boumali, E.O. Silva, Ann. Phys. 458, 169479 (2023)
  • 14 (2025) A. Bouzenada, A. Boumali, R. Vitória, F. Ahmed, M. Al-Raeei, Nucl. Phys. B 994, 116288 (2023)
  • 15 (2025) A. Bouzenada, A. Boumali, F. Ahmed, Nucl. Phys. B 1007, 116682 (2024)
  • 16 (2025) A.Bouzenada and A.Boumali. Ann. Physics 27, 169302 (2023).
  • 17 (2025) E. Schrodinger, Sitz. Preuss. Akad. Wiss. Berlin 105 (1932).

### v2

@@ IV Feshbach-Villars of a particle of spin-1/2 in curved spacetime: The formalism
Now, to simplify the handling of our quantum mechanics problem, let us work in a space of lower dimensions, and because there is no structure in the $z$ direction, we can remove it. Thus, the metric of a static cosmic string with cylindrical symmetry in two dimensions has the following form Vilenkin (1981); Vilenkin and Shellard (1994)
$ds^{2}=dt^{2}-dr^{2}-\alpha^{2}r^{2}d\varphi^{2}$ (72)
where the components of the metric and the inverse metric tensors are, respectively,
$g_{\mu\nu}=\left(\begin{array}[]{ccc}1&0&0\\ 0&-1&0\\ 0&0&-(\alpha r)^{2}\end{array}\right),\quad g^{\mu\nu}=\left(\begin{array}[]{ccc}1&0&0\\ 0&-1&0\\ 0&0&\frac{-1}{(\alpha r)^{2}}\end{array}\right)$ (73)
Note here that, in the Hamiltonian formulation, the evolution parameter is the coordinate time $t$ that labels the foliation by spacelike hypersurfaces $\Sigma_{t}$ defined by $t=\text{const}$. This is the same parameter that appears in the Schrödinger-type equation $i\partial_{t}\Psi=\mathcal{H}\Psi$. For a general \$3+1\$ decomposition of the spacetime metric (ADM formalism), the line element is written as:
$ds^{2}=(N^{2}-h_{ij}N^{i}N^{j})dt^{2}-2h_{ij}N^{j}dx^{i}dt-h_{ij}dx^{i}dx^{j},$ (74)
where $N(t,x)$ is the lapse function and $N^{i}(t,x)$ is the shift vector Arnowitt1962; Wald (1984); Gourgoulhon (2007).
Physically, the lapse function $N$ controls the proper-time separation between neighboring hypersurfaces as measured by Eulerian observers normal to $\Sigma_{t}$. The shift vector $N^{i}$ encodes how the spatial coordinates are shifted when evolving from $\Sigma_{t}$ to $\Sigma_{t+dt}$, representing the gauge freedom in the choice of spatial coordinates. Consequently, the Hamiltonian flow is naturally parametrized by $t$. The relationship between $t$ and the proper time $\tau$ along a given worldline depends on $N$ and $N^{i}$.
For the static diagonal cosmic-string metric considered in Section V, we have $g_{0i}=0$ and $g_{00}=1$. This implies a lapse of $N=\sqrt{g_{00}}=1$ and a vanishing shift vector $N^{i}=0$. Therefore, for static observers (fixed spatial coordinates), the proper time satisfies $d\tau=\sqrt{g_{00}}dt=dt$. In this specific case, the Hamiltonian-flow parameter coincides with both the metric time coordinate and the proper time of static observers.
Now, to evaluate the operators $Y$ and $O$ in the Feshbach-Villars (FV) formalism for a spin-1/2 particle in the static cosmic string spacetime, we start with the metric:
$ds^{2}=dt^{2}-dr^{2}-\alpha^{2}r^{2}d\varphi^{2},$ (75)
with the metric tensor $g_{\mu\nu}=\text{diag}(1,-1,-\alpha^{2}r^{2})$. The inverse metric is $g^{\mu\nu}=\text{diag}(1,-1,-1/(\alpha^{2}r^{2}))$, and $\sqrt{-g}=\alpha r$.
From the definition in the FV formalism:
$Y=\frac{1}{2g_{00}\sqrt{-g}}\{D_{i},\sqrt{-g}g^{0i}\}.$ (76)
Since the metric is diagonal, $g^{0i}=0$ for all spatial indices $i$. This implies:$Y=0.$ The operator $O$ is given by:
$O=\frac{1}{2m}\left[\frac{1}{\sqrt{-g}}D_{i}\left(\sqrt{-g}g^{ij}D_{j}\right)+\frac{ie}{2}\frac{F_{\mu\nu}\sigma^{\mu\nu}}{g_{00}}\right].$ (77)
Assuming no electromagnetic field ($F_{\mu\nu}=0$) and focusing on the gravitational contribution, we simplify:
$O=\frac{1}{2m}\frac{1}{\sqrt{-g}}D_{i}\left(\sqrt{-g}g^{ij}D_{j}\right).$ (78)
In cylindrical coordinates ($r,\varphi$):
$O=\frac{1}{\alpha r}\left[-D_{r}(\alpha rD_{r})-D_{\varphi}\left(\frac{1}{\alpha r}D_{\varphi}\right)\right].$ (79)
Expanding the covariant derivatives ($D_{i}=\partial_{i}+\Gamma_{i}$) and simplifying for the static cosmic string metric (where the spin connection $\Gamma_{i}$ contributes terms proportional to $\alpha$), we obtain:
$O=-\frac{1}{2m}\left[\partial_{r}^{2}+\frac{1}{r}\partial_{r}+\frac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}\right].$ (80)
The FV Hamiltonian for the spin-1/2 particle becomes:
$H=\sigma_{z}\frac{m}{2}+\sigma_{x}O,$ (81)
leading to the coupled system:
$i\partial_{t}\begin{pmatrix}\phi\\ \chi\end{pmatrix}=\left(\frac{m}{2}\sigma_{z}-\frac{1}{2m}\sigma_{x}\left[\partial_{r}^{2}+\frac{1}{r}\partial_{r}+\frac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}\right]\right)\begin{pmatrix}\phi\\ \chi\end{pmatrix}$. (82)
Fron the last equation, The system of equations for $\phi$ and $\chi$ derived from the given Hamiltonian is:
$i\partial_{t}\phi=\dfrac{m}{2}\phi-\dfrac{1}{2m}\left(\partial_{r}^{2}+\dfrac{1}{r}\partial_{r}+\dfrac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}\right)\chi$ (83)
$i\partial_{t}\chi=-\dfrac{m}{2}\chi-\dfrac{1}{2m}\left(\partial_{r}^{2}+\dfrac{1}{r}\partial_{r}+\dfrac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}\right)\phi$ (84)
The FV formalism rewrites the wave function as:
$\psi=\phi+\chi$ (85)
and relates the time derivative to the difference:
$i\partial_{t}\psi=m(\phi-\chi)$ (86)
This allows you to express $\phi$ and $\chi$ in terms of $\psi$ and its time derivative:
$\phi=\frac{1}{2}\left(\psi+\frac{i}{m}\partial_{t}\psi\right)$ (87)
$\chi=\frac{1}{2}\left(\psi-\frac{i}{m}\partial_{t}\psi\right)$ (88)
From some algebraic steps, we obtain a differential equation as:
$\left[\partial_{r}^{2}+\frac{1}{r}\partial_{r}+\frac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}+E^{2}-m^{2}\right]\psi=0$ (89)
To solve the partial differential equation, we employ separation of variables in cylindrical coordinates. Assuming
$\psi(r,\varphi)=R(r)\Phi(\varphi)$ (90)
, the equation splits into angular and radial components. The angular part satisfies:
$\Phi^{\prime\prime}(\varphi)+\alpha^{2}l^{2}\Phi(\varphi)=0,$ (91)
with periodic boundary conditions $\Phi(\varphi+2\pi)=\Phi(\varphi)$. The solution is:
$\Phi(\varphi)=e^{il\varphi},\quad l\in\mathbb{Z}.$ (92)
Substituting $\Phi(\varphi)$ into the original equation yields:
$R^{\prime\prime}(r)+\frac{1}{r}R^{\prime}(r)+\left(k^{2}-\frac{l^{2}}{\alpha^{2}r^{2}}\right)R(r)=0,$ (93)
where $k^{2}=E^{2}-m^{2}$.
This is a Bessel equation of order $\nu=l/\alpha$. The general solution is Andrews et al. (1999); Arfken and Weber (2005):
$R(r)=AJ_{l/\alpha}(kr)+BY_{l/\alpha}(kr),$ (94)
where $J_{\nu}$ and $Y_{\nu}$ are Bessel functions of the first and second kind. For regularity at $r=0$, $B=0$, leaving:
$R(r)=AJ_{l/\alpha}(kr).$ (95)
Combining the radial and angular components, the general solution is:
$\psi(r,\varphi)=\sum_{l=-\infty}^{\infty}A_{l}J_{l/\alpha}(kr)e^{il\varphi}.$ (96)
The energy $E$ is related to the wave number $k$ by:
$E=\pm\sqrt{m^{2}+k^{2}}.$ (97)
Boundary conditions (e.g., $R(R)=0$ at radius $R$) quantize $k$ via zeros of $J_{l/\alpha}(kR)$.
IV.2.1 FV-1/2 oscillator
Now, to include the Dirac oscillator term $im\omega\beta r$ into eq. (2), we proceed with the following substitution in the radial momentum component $\partial_{r}\rightarrow\partial_{r}+m\omega\rho$. Hence, eq. (89) is transformed into
$\left[\left(\partial_{r}-m\omega\rho\right)\left(\partial_{r}+m\omega\rho\right)+\frac{1}{r}\partial_{r}+\frac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}+E^{2}-m^{2}\right]\psi=0$ (98)
or
$\left[\partial_{r}^{2}-m^{2}\omega^{2}\rho^{2}+\frac{1}{r}\partial_{r}+\frac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}+E^{2}-m^{2}+m\omega\right]\psi=0$ (99)
To solve the given partial differential equation (PDE) we employ the separation of variables method, assuming a solution of the form
$\psi(r,\varphi)=R(r)\Phi(\varphi)$ (100)
Separate the angular component by substituting $\psi=R(r)\Phi(\varphi)$ into the PDE and dividing by $R\Phi$:
$\frac{1}{\alpha^{2}r^{2}}\frac{\partial_{\varphi}^{2}\Phi}{\Phi}=-\nu^{2}\quad\Rightarrow\quad\Phi^{\prime\prime}+\alpha^{2}\nu^{2}\Phi=0.$ (101)
The periodic boundary condition $\Phi(\varphi+2\pi)=\Phi(\varphi)$ quantizes $\nu$ as an integer ($\nu\in\mathbb{Z}$). The solution is:
$\Phi(\varphi)=e^{i\nu\alpha\varphi}.$ (102)
Substitute $\Phi(\varphi)$ back into the PDE to obtain the radial equation:
$R^{\prime\prime}+\frac{1}{r}R^{\prime}+\left[E^{2}-m^{2}+m\omega-m^{2}\omega^{2}r^{2}-\frac{\nu^{2}}{\alpha^{2}r^{2}}\right]R=0.$ (103)
This resembles a quantum harmonic oscillator equation with an additional $1/r^{2}$ term Greiner (2001); Flugge (1974) . The Analyze Asymptotic Behavior, we have
  • At $r\to 0$: The dominant term is $\frac{\nu^{2}}{\alpha^{2}r^{2}}$.
  • At $r\to\infty$: The dominant term is $m^{2}\omega^{2}r^{2}$.
This suggests an ansatz:
$R(r)=r^{|\nu|/\alpha}e^{-\frac{1}{2}m\omega r^{2}}F(r)$ (104)
The power of $r$ handles the singularity at the origin, and the exponential ensures normalizability at infinity. Plug $R(r)$ into the radial equation and simplify. You will obtain a differential equation for $F(r)$ of the form:
$F^{\prime\prime}(r)+\left(\frac{2|\nu|/\alpha+1}{r}-2m\omega r\right)F^{\prime}(r)+\left[\kappa-2m\omega\frac{|\nu|}{\alpha}-m\omega\right]F(r)=0,$ (105)
where
$\kappa=E^{2}-m^{2}+m\omega$ (106)
we recognize it as a form of the associated Laguerre equation Greiner (2001); Flugge (1974) .
Let $x=m\omega r^{2}$. Changing variables and simplifying reduces the equation to:
$x\frac{d^{2}F}{dx^{2}}+\left(\frac{|\nu|}{\alpha}+1-x\right)\frac{dF}{dx}+\left[\frac{\kappa-2m\omega|\nu|/\alpha-m\omega}{4m\omega}\right]F=0.$ (107)
This matches the standard associated Laguerre equation:
$xy^{\prime\prime}+(k+1-x)y^{\prime}+ny=0,$ (108)
where: $k=|\nu|/\alpha$, and
$n=\frac{\kappa-2m\omega|\nu|/\alpha-m\omega}{4m\omega}$ (109)
For polynomial solutions (associated Laguerre polynomials $L_{n}^{(k)}(x)$), $n$ must be a non-negative integer.
This imposes:
$\kappa=4m\omega n+2m\omega\frac{|\nu|}{\alpha}+m\omega.$ (110)
Recalling $\kappa=E^{2}-m^{2}+m\omega$, we solve for $E$:
$E^{2}=m^{2}+4m\omega n+2m\omega\frac{|\nu|}{\alpha}.$ (111)
The discrete energy levels are:
$E=\pm\sqrt{m^{2}+2m\omega\left(2n+\frac{|\nu|}{\alpha}\right)},$ (112)
where $n=0,1,2,\ldots$ is the radial quantum number, and $\nu\in\mathbb{Z}$ is the angular quantum number.
The full solution combines radial and angular components:
$\psi(r,\varphi)=\sum_{n,\nu}C_{n,\nu}\,r^{|\nu|/\alpha}e^{-m\omega r^{2}/2}L_{n}^{|\nu|/\alpha}(m\omega r^{2})\,e^{i\nu\alpha\varphi},$ (113)
where $C_{n,\nu}$ are normalization constants. The energy spectrum is discrete and depends on $n$, $\nu$, and $\alpha$.
V FV oscillator in a spinning cosmic string with torsion and dislocation
In this third example we consider a spinning cosmic string carrying both a time–like torsion (frame–dragging) and a space–like torsion (screw dislocation). A convenient form of the corresponding line element (in cylindrical coordinates) is Puntigam and Soleng (1997); Ozdemir (2005); Jusufi (2016)
$ds^{2}=\left(dt+a\,d\varphi\right)^{2}-dr^{2}-\alpha^{2}r^{2}d\varphi^{2}-\left(dz+\beta\,d\varphi\right)^{2},$ (114)
where $\alpha\in(0,1]$ is the angular–deficit parameter, $a$ is proportional to the linear density of angular momentum (time–like torsion) and $\beta$ encodes the screw–dislocation (space–like torsion). Note that $g_{\varphi\varphi}=a^{2}-\alpha^{2}r^{2}-\beta^{2}$; in the regime $a^{2}>\beta^{2}$ one must restrict the radial coordinate to $r>r_{0}=\sqrt{a^{2}-\beta^{2}}/\alpha$ in order to avoid unphysical closed time–like curves, so that $r=r_{0}$ plays the role of a geometric hard wall Garah and Boumali (2025b).
A convenient orthonormal coframe $\{\hat{\theta}^{a}\}$ adapted to the line element (114) is
$\hat{\theta}^{0}=dt+a\,d\varphi,\qquad\hat{\theta}^{1}=dr,\qquad\hat{\theta}^{2}=\alpha r\,d\varphi,\qquad\hat{\theta}^{3}=dz+\beta\,d\varphi,$ (115)
so that $ds^{2}=\eta_{ab}\,\hat{\theta}^{a}\hat{\theta}^{b}$ with $\eta_{ab}=\mathrm{diag}(1,-1,-1,-1)$. The corresponding tetrad $e^{a}{}_{\mu}$ and its inverse $e_{a}{}^{\mu}$ (defined by $\hat{\theta}^{a}=e^{a}{}_{\mu}\,dx^{\mu}$ and $dx^{\mu}=e_{a}{}^{\mu}\,\hat{\theta}^{a}$) read
$\left(e^{a}{}_{\mu}\right)=\begin{pmatrix}1&0&a&0\\ 0&1&0&0\\ 0&0&\alpha r&0\\ 0&0&\beta&1\end{pmatrix},\qquad\left(e_{a}{}^{\mu}\right)=\begin{pmatrix}1&0&-\dfrac{a}{\alpha r}&0\\ 0&1&0&0\\ 0&0&\dfrac{1}{\alpha r}&0\\ 0&0&-\dfrac{\beta}{\alpha r}&1\end{pmatrix}$, (116)
where the coordinate ordering is $(t,r,\varphi,z)$ and the local Lorentz index ordering is $(0,1,2,3)$.
In the notation of Sec. II, the curved-space Dirac matrices are $\gamma^{\mu}=e_{a}{}^{\mu}\gamma^{a}$, where $\gamma^{a}$ satisfy $\{\gamma^{a},\gamma^{b}\}=2\eta^{ab}$. In the standard Dirac representation,
$\gamma^{0}=\begin{pmatrix}\mathbb{I}_{2}&0\\ 0&-\mathbb{I}_{2}\end{pmatrix},\qquad\gamma^{i}=\begin{pmatrix}0&\sigma^{i}\\ -\sigma^{i}&0\end{pmatrix}\ (i=1,2,3),\qquad\alpha^{i}=\gamma^{0}\gamma^{i},\ \beta=\gamma^{0},$ (117)
with $\sigma^{i}$ the Pauli matrices and $\mathbb{I}_{2}$ the $2\times 2$ identity. Using (116), the explicit curved matrices needed in the Hamiltonian/FV form are
$\gamma^{t}=\gamma^{0}-\frac{a}{\alpha r}\gamma^{2},\qquad\gamma^{r}=\gamma^{1},\qquad\gamma^{\varphi}=\frac{1}{\alpha r}\gamma^{2},\qquad\gamma^{z}=\gamma^{3}-\frac{\beta}{\alpha r}\gamma^{2}.$ (118)
For $r\neq 0$, the only $r$-dependent coframe component is $\hat{\theta}^{2}=\alpha r\,d\varphi$, and the torsion is distributional at the string core (screw dislocation) while vanishing in the exterior region. In the exterior region one finds the single nonvanishing Lorentz connection component
$\omega_{\varphi}{}^{12}=1-\alpha,\qquad\Rightarrow\qquad\Gamma_{\varphi}=\frac{1}{4}\,\omega_{\varphi}{}^{12}\,\gamma_{12}=\frac{1-\alpha}{4}\,\gamma^{1}\gamma^{2},$ (119)
where $\Gamma_{\mu}=\frac{1}{4}\omega_{\mu}{}^{ab}\gamma_{ab}$ is the spinor connection used in Sec. II and $\gamma_{ab}=\frac{1}{2}[\gamma_{a},\gamma_{b}]$. The parameter $\alpha\in(0,1]$ encodes the angular deficit (for a straight string $\alpha=1-4\mu$), $a$ is proportional to the angular momentum per unit length (“spinning” string), and $\beta$ is the screw-dislocation/torsion parameter (Burgers vector per $2\pi$) that shifts the identification along $z$ under $\varphi\to\varphi+2\pi$. In the FV–oscillator problem, $m$ and $\omega$ denote the particle mass and the oscillator frequency, while $l\in\mathbb{Z}$ and $k\in\mathbb{R}$ are the azimuthal and longitudinal quantum numbers, respectively.
Because the background is stationary and invariant under translations along $z$, we separate variables as
$\psi\left(t,r,\varphi,z\right)=e^{-iEt}\,e^{il\varphi}\,e^{ikz}\,R(r),$ (120)
with $l\in\mathbb{Z}$ and $k\in\mathbb{R}$. The mixed components $g_{t\varphi}$ and $g_{z\varphi}$ imply that the azimuthal dependence enters through the combination $(\partial_{\varphi}+a\,\partial_{t}-\beta\,\partial_{z})$ . Consequently, the effective angular index is shifted to
$\nu_{\mathrm{eff}}=l+aE-\beta k.$ (121)
Following the same FV–oscillator steps as in the previous subsection, we obtain the radial equation
$\frac{d^{2}R}{dr^{2}}+\frac{1}{r}\frac{dR}{dr}+\left[\left(E^{2}-m^{2}-k^{2}\right)-m^{2}\omega^{2}r^{2}-\frac{\nu_{\mathrm{eff}}^{2}}{\alpha^{2}r^{2}}\right]R=0.$ (122)
Upon introducing the variable $x=m\omega r^{2}$, the regularity at the origin and normalizability yield the quantization condition
$E^{2}=m^{2}+k^{2}+4m\omega n+\frac{2m\omega}{\alpha}\left|\nu_{\mathrm{eff}}\right|,\qquad n=0,1,2,\ldots$ (123)
which is implicit in the energy through $\nu_{\mathrm{eff}}$ (a characteristic feature of spinning backgrounds). For the branch $\nu_{\mathrm{eff}}>0$, Eq. (123) becomes a quadratic equation for $E$ and can be written explicitly as
$E=\frac{m\omega a}{\alpha}\pm\sqrt{\left(\frac{m\omega a}{\alpha}\right)^{2}+m^{2}+k^{2}+4m\omega n+\frac{2m\omega}{\alpha}\left(l-\beta k\right)}.$ (124)
For $\nu_{\mathrm{eff}}<0$ we obtain instead
$E=-\frac{m\omega a}{\alpha}\pm\sqrt{\left(\frac{m\omega a}{\alpha}\right)^{2}+m^{2}+k^{2}+4m\omega n-\frac{2m\omega}{\alpha}\left(l-\beta k\right)}.$ (125)
In the limits $a\rightarrow 0$ and $\beta\rightarrow 0$ we recover the static cosmic string result of Eq. (111); the flat–space Moshinsky spectrum follows further for $\alpha\rightarrow 1$. This provides a nontrivial consistency check of the FV construction: in all limits where the spectrum is known from the standard Dirac analysis, the observable eigenvalues coincide, as expected for a Hamiltonian reformulation.
VI Conclusion
This work extends the Feshbach–Villars (FV) formalism to spin–$1/2$ particles in curved spacetime by deriving a Hamiltonian representation of the covariant Dirac dynamics and implementing a diagonal FV transformation. The resulting two-component framework makes the separation between positive- and negative-frequency sectors explicit and provides a compact way to track gravitational and electromagnetic couplings, including spin–field interaction terms, in both $(1+2)$ and $(1+3)$ dimensions.
Discussion. The applications to static cosmic-string geometries (with and without oscillator couplings) illustrate how geometric parameters (such as the deficit angle and torsion/dislocation data) control the quantization conditions, spectra, and wave-function profiles. In this sense, the FV representation serves as a practical bridge between standard Dirac dynamics and second-order formulations in which particle–antiparticle structure is less transparent.
In addition, we have made explicit two points that are sometimes only implicit in second-order approaches: (i) the FV Hamiltonian is pseudo-Hermitian with respect to a simple metric operator $\eta_{FV}=\sigma_{3}\otimes\mathbb{I}_{4}$, so that the conserved quantity is a charge-type norm, and (ii) the equivalence with the standard Dirac equation is understood as a restriction to the physical (Dirac) subspace that removes the spurious solutions of the squared equation. Finally, the role of stationarity is particularly transparent: when $g_{0i}\neq 0$ (nonvanishing ADM shift), the FV operator $\mathrm{Y}$ does not vanish and this directly feeds into the energy quantization. The spinning string example exhibits this through the effective angular index $\nu_{\mathrm{eff}}=l+aE-\beta k$ and the analytic spectra in Eqs. (124)–(125).
Limitations. The present approach is a single-particle relativistic quantum-mechanical description on a prescribed background. It does not include intrinsically field-theoretic effects in curved spacetime, such as particle creation, vacuum polarization, or gravitational backreaction, and it does not resolve the well-known ambiguity of particle concepts in generic nonstationary geometries. Moreover, global issues (e.g., the existence of a suitable foliation and boundary conditions ensuring self-adjointness in the FV inner product) can restrict the applicability of a Hamiltonian treatment. These limitations help explain why semiclassical and quantum-field-theoretic methods have long been central to the study of curved-spacetime phenomena.
Within its domain of validity, the FV formulation developed here provides a systematic platform for further studies, including extensions to broader classes of stationary metrics, controlled comparisons with semiclassical/QFT treatments, and investigations of spinor dynamics in other topologically nontrivial backgrounds.
VII Data Availability Statement
No Data associated in the manuscript